lima corp has a wacc of 10%. The company cost of equity is 13% and its cost of debt is 7%. If tax rate is 35%, what is firms debt equity ratio?
Answer #1

The weighted average cost of capital is calculated using the following equation

WACC = Wd rd (1 - tc ) + We re + Wp rp

Weight of equity can be be written as (1 - Wd ) since We + Wd = 1

WACC = Wd rd (1 - tc ) + ( 1 - Wd ) re

0.10 = Wd 0.07 ( 1 - 0.35) + (1 - Wd ) 0.13

0.10 = 0.0455 Wd + ( 1 - Wd ) 0.13

0.10 = 0.0455 Wd + 0.13 - 0.13 Wd

-0.03 = -0.0845 Wd

Wd = 35.50%

Weight of debt = 35.50%

Weight of equity = 1 - 0.3550 = 0.6450

Weight of equity = 64.50%